OpenStudy (anonymous):

how to multiply a whole number with a fraction for example 3 X 1/5 i need steps

hero (hero):

Multiply 3 times 5, then add one. Afterwards put the number you get over 5.

OpenStudy (anonymous):

3/1x1/5 would equal 3/5 Multiply numerators together, then multiply denominators together. To turn 3 into a fraction, put it over 1. 3/1=3. So then multiply 3 by 1 and 1 by 5 to get 3/5

hero (hero):

\[3 \frac{1}{5}\] That's what it is. Then after doing what I suggested you end up with: \[\frac{16}{5}\]

OpenStudy (anonymous):

thanks i get it for this problem but multiplying fractions has its rule for every different problem

OpenStudy (debbieg):

I think she is multiplying whole # x fraction, not converting a mixed number to an improper fraction...?

hero (hero):

Oh I see how that can be confusing.

hero (hero):

Shameful. They should have kept mixed fractions in this form: \[3 + \frac{1}{5}\] Which would be distinguishable from \[3 \dot\ \frac{1}{5}\]

OpenStudy (debbieg):

Like @Deanozaur said, it's sometimes easier to do this problem type if you write the whole number so that you can "treat it" like a fraction: \(5\cdot \dfrac{1}{3}=\dfrac{5}{1}\cdot \dfrac{1}{3}\) and then just multiply across num'r, and across den'r, as with any fraction multiplication.

OpenStudy (debbieg):

\(5\cdot \dfrac{1}{3}=\dfrac{5}{1}\cdot \dfrac{1}{3}= \dfrac{5\cdot1}{1\cdot3}= \dfrac{5}{3}\)

hero (hero):

Yes, but @DebbieG I can see how one could get confused if they saw a mixed fraction of this form: \[3 \frac{1}{5}\] That's all I was pointing out

hero (hero):

That also looks like 3 times 1/5 even though it isn't

OpenStudy (debbieg):

The main point to remember is that you write any whole number as a fraction, just by putting it over a den'r of 1.

OpenStudy (debbieg):

Huh? @Hero I was addressing the op's question. I wasn't responding to you, other than the one post where I pointed out that you had misinterpreted the question.

OpenStudy (debbieg):

I do see your point, although I thought her question "how to multiply a whole number with a fraction for example 3 X 1/5" was pretty clear what she meant....

hero (hero):

I thought she was referring to the process of multiplying with mixed numbers.

OpenStudy (anonymous):

Thanks for the medal :)

OpenStudy (debbieg):

Oh, I did 5 x 1/3, not 3 x 1/5, but the concept is the same so you can see the example I did, @suzy4321

hero (hero):

See @DebbieG, we all make mistakes.

hero (hero):

I thought she was referring to the process of multiplying with mixed numbers.

hero (hero):

That also looks like 3 times 1/5 even though it isn't

hero (hero):

Yes, but @DebbieG I can see how one could get confused if they saw a mixed fraction of this form: \[3 \frac{1}{5}\] That's all I was pointing out

OpenStudy (debbieg):

Yes, we all make mistakes. I wasn't putting you down for your mistake; I was pointing it out so that her ACTUAL question could be answered properly. Yes, I know what you thought she was referring to. Yes, I suppose that a person could mix up mixed numerals and the product of a whole number and a fraction, although the notation is not identical, so it's really just a matter of not using sloppy notation. I was just trying to help her with her question, did not mean to start a debate with you.

hero (hero):

Why do you have to assume that we're having a debate rather than a discussion?